OpenStudy (anonymous):

"I am seven times as old as you were when i was as old as you are" , said a man to his son. Find their present ages.if the sum of their ages is 110 years?

OpenStudy (anonymous):

Let their ages be x and y The sum is 110 so x+y=110

OpenStudy (anonymous):

ok then?

OpenStudy (anonymous):

that's not right

OpenStudy (mathstudent55):

Let the present father's age be f. Let the present son's age be s. f + s = 110 When the father was as old as the son is now, it was f - s years ago. The son was then s - (f - s) = s - f + s = 2s - f years old f = 7(2s - f) f = 14s - 7f 8f = 14s f = (14/8) s f + s = 110 (14/8)s + s = 110 s = 40 f = 70 The father is 70 and the son is 40. 30 years ago, the father was 40 and the son was 10. The father is now 7 times older than when the son was 10, which is when the father was the son's current age of 40.

OpenStudy (anonymous):

ok i got it..thnxs

OpenStudy (mathstudent55):

The son is s years old now, and the father is f years old now. When the father was the age the son is now, that was f - s years ago.. Ages Now Old Time was Ages in the Past Father f f - s years ago f - (f - s) Son s f - s years ago s - (f - s) "I am seven times as old as you were when i was as old as you are" "I am" means the father's age now, f. "seven times as old" means 7 * " as you were " that means s - (f - s) "when i was as old as you are" This translates into the equation f = 7[s - (f - s)] Which can be simplified to f = 7(s - f + s) f = 7(2s - f) f = 14s - 7f